OpenStudy (anonymous):

What is an example of a function that can only be differentiated to the third degree @Mathematics

OpenStudy (anonymous):

\[y = 3x{^3} - 5x{^2} + 3x - 4\]\[y' = 9x{^2} - 10x + 3\]\[y'' = 18x - 10\]\[y'''=18\]\[y''''=0\]

OpenStudy (anonymous):

What is your reasoning?

OpenStudy (anonymous):

I thought polynomials are always infinitely differentiable?

OpenStudy (anonymous):

oh i went into flow only 3. so sorry my example will go upto to fifth differential